OpenStudy (anonymous):

-6+5b=6b-7 what is the solution of this equation?

OpenStudy (anonymous):

b=1

OpenStudy (jamesj):

b = 1. Put it in the equation and see!

OpenStudy (anonymous):

\[-6+5b=6b-7\] \[5b-6b=-7+6\] \[-b=-1\] \[-1(b) = -1(1)\] \[b=1\]

OpenStudy (anonymous):

It's right